OpenStudy (anonymous):

factor by any method 2x^4-21x^2-11

OpenStudy (anonymous):

factor \[2z^4-21z-11\] then replace \(z\) by \(x^2\)

OpenStudy (anonymous):

ok then what is next

OpenStudy (anonymous):

oh damn i made a typo, should have written \[2z^2-21z-11\]

OpenStudy (compassionate):

You can start by trying to put them into two separate binomials. @satellite73 , also, please check the reports or view [ http://openstudy.com/study#/updates/52ec7367e4b0a362a894b4ea ]

OpenStudy (triciaal):

let x^2 = b then what you have that is more recognizable is 2b^2-21b-11 to factor, multiply the coefficient of b^2 by the constant -11 to get -22 what are the factors of -22 such that when you add them you get -21 yes the factors would be -22, +1 so, 2b^2 -22b +1 b -11 regroup and factor -11(2b +1) +b(2b +1) (2b +1)(b-11) remember b= x^2 so (2x^2 + 1)(x^2 + 1)